OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (anonymous):

Any line perpendicular to a line described by -7x + 9y = C will have the equation: 9x+7y = K If the line passes through (-7, 6), then: K = 9(-7) + 7(6) = -63 + 42 = -21 The line (standard form) is 9x+7y=-21 in slope intercept form: y = -9x/7 - 3 source: http://answers.yahoo.com/question/index?qid=20110802192756AAhrGSP

OpenStudy (anonymous):

Thank you(:

OpenStudy (anonymous):

all is well !!!:)